a balloon contains 1.82 mol of oxygen gas, O2(g), at 303K and a pressure of 452 Torr. what is the volume, in liters, of the balloon?
1
Expert's answer
2020-03-23T10:14:25-0400

P= 452 Torr = 0,595 atm ; R = 0,082 ; n = 1,82 mol ; T = 303 K ;

V=nRT/P=1.82*0.082*303/0.595=76(L)